Use the given degree of confidence and sample data to construct a confidence interval for the population proportion p.


16)n = 182, x = 135; 95 percent


Answer by stanbon(75887)

Use the given degree of confidence and sample data to construct a confidence interval for the population proportion p.
---------------
16)n = 182, x = 135; 95 percent
-----------------
sample proportion: p-hat = 135/182 = 0.74
----
E = 1.96*sqrt[0.74*0.26/182] = 0.0637
------------
95% CI: 0.74-0.0637 < p < 0.74+0.0637
=========================================
Cheers,
Stan H.
================